# Instant Runoff Voting and the Reinforcement Paradox

### v2

@@ Abstract.
We analyze the susceptibility of instant runoff voting (IRV) to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two sets of ballots such that a given losing candidate wins each of the sub-elections. Using these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We also analyze the frequency with which the paradox occurs in a large dataset of real-world ranked-choice elections to provide empirical probabilities. Our general finding is that IRV is highly susceptible to this paradox in three-candidate elections.
@@ 1. Introduction
The voting method of instant runoff voting (IRV), also referred to as the alternative vote, the plurality elimination rule, and colloquially as “ranked-choice voting” in the United States, is famously susceptible to many classical voting paradoxes [14, 32, 41]. In this article we analyze the susceptibility of IRV to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two subsets of ballots such that a given losing candidate wins each of the sub-elections. Using these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We then analyze the frequency of the paradox in a large dataset of real-world ranked-choice elections to provide empirical probabilities. By characterizing the three-candidate IRV elections which demonstrate the reinforcement paradox and estimating its frequency, we view our work as a contribution to the larger agenda laid out by Felsenthal in [12].
@@
Reinforcement paradoxes have received attention in the social choice literature due to their interesting and surprising nature, much like other paradoxes to which IRV is susceptible, such as upward or downward monotonicity paradoxes. However, unlike monotonicity paradoxes, IRV’s susceptibility to reinforcement paradoxes has direct administrative implications: all ballots must be transmitted—either electronically or physically—to a central counting location. In contrast, a simpler system like plurality allows each precinct to report vote totals individually, and a winner can be determined by summing these across precincts. Due to IRV’s vulnerability to reinforcement paradoxes, precinct-by-precinct counting is not possible. As a result, IRV ballot tabulation may experience administrative delays that would not occur with methods like plurality. Thus, a study of reinforcement paradoxes is motivated by practical as well as theoretical considerations.
@@
Our work in this paper connects to three previous strands of the literature concerning IRV and voting paradoxes. The first gives conditions on a preference profile which characterize when an election demonstrates a given paradox. Examples are Lepelley et al. [21], Miller [29], and Ornstein and Norman [33], which characterize the three-candidate profiles which demonstrate an upward or downward monotonicity paradox under IRV. Similarly, Graham-Squire [16] characterizes three-candidate profiles which demonstrate a no-show paradox under IRV, including profiles which contain partial ballots. Kamwa et al. [18] study a variety of voting rules, but in particular provide conditions under which three-candidate profiles produce paradoxes such as negative participation paradoxes under IRV.
@@
The second strand of literature typically uses the conditions given by the works cited above to estimate the frequency with which a given paradox occurs under IRV. For example, Ornstein and Norman [33] estimate the frequency of monotonicity failures in IRV elections when the ballot data is generated using a spatial model. Similarly, Lepelley et al. [21] and Miller [29] estimate the frequency with which IRV produces monotonicity paradoxes under various theoretical models of voter behavior. Lepelley and Merlin [22] analyze the frequency with which IRV demonstrates a no-show paradox in three-candidate elections under the impartial culture and impartial anonymous culture models. Kamwa et al. [18] analyze the probability that IRV (among other methods) demonstrates no-show or negative participation paradoxes on single-peaked domains.
@@
Previous work on the reinforcement paradox is fairly limited. The notion of a reinforcement paradox dates to the work of Young [45, 46], who defined the concept and proved essentially that the only reasonable voting methods not susceptible to the paradox are positional scoring rules. Felsenthal and Nurmi [13] show that various Condorcet inconsistent voting rules (including IRV) are susceptible to the paradox even in “tame” profiles in which a Condorcet winner exists and the voting rule chooses this candidate. Brandt et al. [2] provide some “minimal” examples of the reinforcement paradox under IRV where the number of candidates and voters is small. Brandt et al. [4] analyze the frequency of reinforcement paradoxes under various Condorcet consistent voting rules. We are aware of only two previous papers which attempt to measure the frequency with which IRV produces reinforcement paradoxes, both of which focus on the three-candidate case. Plassmann and Tideman [37] measure this frequency under a spatial model, and Courtin et al. [8] use the impartial culture and impartial anonymous culture models of voter behavior. Our work differs from these studies in significant aspects which we explicate more fully below. In brief, these two studies tend to generate two elections and then say that the reinforcement paradox has occurred if these two elections have the same winner but when the ballots from the two elections are combined the winner is different. By contrast, we consider only one election and say the paradox has occurred in this election if there exists a way to partition the ballots into two sub-elections where the winners of the sub-elections are the same but differ from the winner of the original election.
@@ 2. Preliminaries
In this section we define IRV and types of reinforcement paradoxes, accompanied by examples.
In an election which uses IRV voters cast preference ballots with a (possibly partial) linear ranking of the candidates. After an election, the ballots are aggregated into a preference profile, which shows the number of each type of ballot cast. For example, Table 1 shows a preference profile involving the three candidates $C$, $G$, and $H$. The number 2211 denotes that 2211 voters ranked $G$ first, $H$ second, and $C$ third. In the real-world IRV elections we analyze in Section 6, voters are allowed to provide partial rankings. Table 1 has three columns where voters do not provide a complete ranking–2003 voters rank only $G$ on their ballots, for example.
An IRV election proceeds in rounds where, in each round, the number of first-place votes for each candidate is calculated. If a candidate has a majority of the (remaining) first-place votes, they are declared the winner. If no candidate receives a majority of votes, the candidate with the fewest first-place votes is eliminated and the names of the remaining candidates on the affected ballots are moved up. The process repeats until a candidate is declared the winner. [^1] The issue of ties does not arise in our work in this paper, and thus we ignore the possibility of multiple winners. In the case when partial ballots are allowed, any ballot in which all candidates have been eliminated is considered “exhausted” and plays no further role in the determination of the winner. [^2] In theory there are many ways we could treat exhausted ballots besides removing them from the election. However, as far as we are aware, any actual election office treats exhausted ballots in this fashion.
@@
Initially, the vote totals for Cacy, Giessel, and Holland are 5521, 5653, and 5535, respectively. Cacy has the fewest votes and is eliminated from the election. As a result, 2230 votes are transferred to Giessel and 418 to Holland, and Giessel wins in the last round 7883 votes to 5953.
@@
Table 2 provides a second partition of the ballots, Partition 2, in which $C$ wins under IRV when using only the ballots from $\mathcal{B}_{1}$ and only the ballots from $\mathcal{B}_{2}$, yet when we merge the ballots to recover the original election data, $C$ loses.
For both Cacy and Holland, in this election win $+$ win $=$ loss in some sense.
@@
Example 2 demonstrates the notion of a reinforcement paradox, the subject of this article. The study of this paradox dates to the work of Young [45]. Young considered a voting method as a function $f$ which takes as input a set of ranked ballots and outputs an election winner (or a set of winners, if there is some sort of tie). If $\mathcal{B}$ is a set of ballots then $f(\mathcal{B})$ denotes the election winner under a voting method $f$, which leads to the definition of consistency found in [45].
@@
A voting method that is not consistent is said to be susceptible to the reinforcement paradox. This paradox has other names, such as the “multiple districts paradox” [14], and is a social choice analogue of the classical Simpson’s Paradox from statistics.
@@
An election with ballot set $\mathcal{B}$ demonstrates a reinforcement paradox under a voting method $f$ if there exists a partition of the ballot set $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ and $f(\mathcal{B}_{1})\neq f(\mathcal{B}_{1}\cup\mathcal{B}_{2})$.
In [46], Young showed that consistency and a few other uncontroversial axioms completely characterize positional scoring rules. Thus, any “reasonable” voting method which is not a scoring rule can demonstrate a reinforcement paradox for some election. We analyze only IRV, and thus the function $f$ represents that voting method throughout the article.
We conclude this section by providing additional notation for preference profiles and giving vocabulary for different types of reinforcement paradox.
Assume the preference profiles for complete and partial rankings are as shown in Tables 3 and 4 where in each case the numbers $X_{i},Y_{i}$, and $Z_{i}$ indicate the number of voters with the corresponding ranking and that these numbers sum to the total number of voters, which we denote $V.$ Note in Table 4 we do not list ballots of length two, as such ballots convey the same ranking information as the corresponding complete ballots.
@@
1st choice $A$ $A$ $B$ $B$ $C$ $C$
2nd choice $B$ $C$ $A$ $C$ $A$ $B$
3rd choice $C$ $B$ $C$ $A$ $B$ $A$
@@
1st choice $A$ $A$ $A$ $B$ $B$ $B$ $C$ $C$ $C$
2nd choice $B$ $C$ $A$ $C$ $A$ $B$
3rd choice $C$ $B$ $C$ $A$ $B$ $A$
@@
When there are three candidates, we refer to the candidate with the smallest number of first-place votes as the plurality loser. This candidate can never be the winner in an IRV election since they are eliminated in the first round. We refer to the remaining losing candidate, (who may receive either the largest or second-largest number of first-place votes) as the runner-up.
Definition 3.
An election demonstrates a plurality loser reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the plurality loser is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ or only the ballots in $\mathcal{B}_{2}$.
Definition 4.
An election demonstrates a runner-up reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the runner-up is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ or only the ballots in $\mathcal{B}_{2}$.
If an election demonstrates both a plurality loser and runner-up paradox, we say the election demonstrates a double reinforcement paradox.
We now turn to the problem of determining if a ballot set $\mathcal{B}$ in a three-candidate election can be partitioned to produce a runner-up or plurality loser reinforcement paradox.
3. Necessary and Sufficient Conditions
The purpose of this article is to investigate the probability with which a three-candidate election demonstrates a reinforcement paradox. To examine such probabilities, we first must know when an election demonstrates a paradox. This section characterizes when an election demonstrates a reinforcement paradox, identifying sets of inequalities in terms of the parameters provided in Table 3 for the case of complete ballots or the parameters provided in Table 4 for the partial ballot case.
First, we observe that if one of the candidates earns an initial majority of first-place votes (we term such a candidate a “majority candidate”) then the election cannot demonstrate a reinforcement paradox. That is, we cannot observe a paradox in the complete ballot case if $X_{1}+X_{2}>Y_{1}+Y_{2}+Z_{1}+Z_{2}$ or $Y_{1}+Y_{2}>X_{1}+X_{2}+Z_{1}+Z_{2}$ or $Z_{1}+Z_{2}>X_{1}+X_{2}+Y_{1}+Y_{2}$; similarly for the partial ballot case. The reason is that if a majority candidate exists, that candidate will earn a majority of first-place votes in either $\mathcal{B}_{1}$ or $\mathcal{B}_{2}$ for any partition of the ballot set.
@@ 3. Necessary and Sufficient Conditions
If an election contains a majority candidate then the election cannot demonstrate a reinforcement paradox.
In the complete ballot case, let $X=X_{1}+X_{2}$, $Y=Y_{1}+Y_{2}$, and $Z=Z_{1}+Z_{2}$; in the partial ballot case, let $X^{\prime}=X_{1}+X_{2}+X_{3}$, $Y^{\prime}=Y_{1}+Y_{2}+Y_{3}$, and $Z^{\prime}=Z_{1}+Z_{2}+Z_{3}$. We assume without loss of generality that $A$ is the IRV winner and $C$ is the plurality loser. Thus, in the complete ballot case we assume
$Z<X,\ Z<Y\text{ and }Y+Z_{2}<X+Z_{1}.$
In the partial ballot case we assume
@@
Under these assumptions, the following propositions characterize when a three-candidate election demonstrates a reinforcement paradox of either type.

### v3

@@ Abstract.
We analyze the susceptibility of instant runoff voting (IRV) to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two sets of ballots such that a given losing candidate wins each of the sub-elections. Applying these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We also analyze the frequency with which the paradox occurs in a large dataset of real-world ranked-choice elections to provide empirical probabilities. Our general finding is that IRV is highly susceptible to this paradox in three-candidate elections.
@@
The voting method of instant runoff voting (IRV), also referred to as the alternative vote, the plurality elimination rule, and colloquially as “ranked-choice voting” in the United States, is famously susceptible to many classical voting paradoxes [22, 42, 52]. In this article we analyze the susceptibility of IRV to a lesser-studied paradox known as a reinforcement paradox, which occurs when candidate $X$ wins under IRV in two distinct elections but $X$ loses in the combined election formed by merging the ballots from the two elections. For three-candidate IRV elections we provide necessary and sufficient conditions under which there exists a partition of the ballot set into two subsets of ballots such that a given losing candidate wins each of the sub-elections. Applying these conditions, we use Monte Carlo simulations to estimate the frequency with which such partitions exist under various models of voter behavior. We then analyze the frequency of the paradox in a large dataset of real-world ranked-choice elections to provide empirical probabilities. By characterizing the three-candidate IRV elections that demonstrate susceptibility to the reinforcement paradox and estimating its frequency, we view our work as a contribution to the larger agenda laid out by Felsenthal in [20].
@@
Reinforcement paradoxes have received attention in the social choice literature due to their interesting and surprising nature, much like other paradoxes to which IRV is susceptible, such as upward or downward monotonicity paradoxes. However, unlike monotonicity paradoxes, IRV’s susceptibility to reinforcement paradoxes has direct administrative implications: to determine the IRV winner, it is not sufficient to know the winner at each precinct. Instead, complete information about the numbers of ballots with each different ranking must be transmitted—either electronically or physically—to a central counting location and collated before the IRV winner can be determined. In contrast, a voting method such as plurality (or a scoring rule) allows each precinct to simply report the number of first-place (and possible second-place, etc.) votes for each candidate, and for the winner to be determined by summing these results across precincts. As a result, IRV ballot tabulation and the subsequent determination of the winner is of necessity more complicated than with methods like plurality. Thus, a study of reinforcement paradoxes is motivated by practical as well as theoretical considerations.
@@
Our work in this paper connects to three previous strands of the literature concerning IRV and voting paradoxes. The first gives conditions on a preference profile which characterize when an election is susceptible to a given paradox. Examples are Lepelley et al. [31], Miller [39], and Ornstein and Norman [43], which characterize the three-candidate profiles which demonstrate an upward or downward monotonicity paradox under IRV. Similarly, Graham-Squire [24] characterizes three-candidate profiles which demonstrate a no-show paradox under IRV, including profiles which contain partial ballots. Kamwa et al. [27] study a variety of voting rules, but in particular provide conditions under which three-candidate profiles produce paradoxes such as negative participation paradoxes under IRV.
@@
The second strand of literature typically uses the conditions given by the works cited above to estimate the frequency with which a given paradox occurs under IRV. For example, Ornstein and Norman [43] estimate the frequency of monotonicity failures in IRV elections when the ballot data is generated using a spatial model. Similarly, Lepelley et al. [31] and Miller [39] estimate the frequency with which IRV produces monotonicity paradoxes under various theoretical models of voter behavior. Lepelley and Merlin [32] analyze the frequency with which IRV demonstrates a no-show paradox in three-candidate elections under the impartial culture and impartial anonymous culture models. Kamwa et al. [27] analyze the probability that IRV (among other methods) demonstrates no-show or negative participation paradoxes on single-peaked domains. McCune and Wilson [38] examine the susceptibility of IRV to negative participation paradoxes.
More broadly, there is a large literature examining the extent to which various voting rules are susceptible to strategic manipulation. Such work need not rely on algebraic characterizations involving the vote totals in the preference profile; see [1, 19, 26, 28] for such examples.
@@
Previous work on the reinforcement paradox is fairly limited. The notion of a reinforcement paradox dates to the work of Young [57, 58], who defined the concept and proved essentially that the only reasonable voting methods not susceptible to the paradox are positional scoring rules. Felsenthal and Nurmi [21] show that various Condorcet inconsistent voting rules (including IRV) are susceptible to the paradox even in “tame” profiles in which a Condorcet winner exists and the voting rule chooses this candidate. Brandt et al. [5] provide some “minimal” examples of the reinforcement paradox under IRV where the number of candidates and voters is small. Brandt et al. [8] analyze the frequency of reinforcement paradoxes under various Condorcet consistent voting rules. We are aware of only two previous papers which attempt to measure the frequency with which IRV produces reinforcement paradoxes, both of which focus on the three-candidate case. Plassmann and Tideman [47] measure this frequency under a spatial model, and Courtin et al. [12] study the paradox under the impartial culture and impartial anonymous culture models of voter behavior. Our work differs from these studies in significant aspects which we explicate more fully below. In brief, the approach taken by these studies is to generate two elections and then say that the reinforcement paradox has occurred if the two elections have the same winner, but when the ballots from the two elections are combined, the winner is different. By contrast, we consider only one election and say that the paradox has occurred in this election if there exists a way to partition the ballots into two sub-elections where the winners of the sub-elections are the same but differ from the winner of the original election. Our work is complementary to theirs and allows us to focus on one set of ballots and the characteristics required for it to demonstrate a susceptibility to this kind of inconsistency.
Another reason for our interest in this approach is that it parallels a well-established line of research regarding self-consistency in allocation rules. For instance, Aumann and Maschler [2] identify consistency (in which the outcome of dividing assets among a set of agents doesn’t change when applied to a subset of agents) as a crucial characteristic of claims methods. Balinski and Young [3] define an analogous notion of uniformity to characterize divisor methods of apportionment. Converse consistency and other related ideas also appear in the cooperative game and bargaining literatures; see [55] for a survey.
Consistency in allocation rules is inherently different from consistency in voting rules: consistency in allocation rules requires that if an allocation is recalculated after a subset of agents “departs” with their share, the allocation does not change. Put another way, if the agents are partitioned into two subsets and the shared goods in each subset are pooled and then reallocated, the results are unchanged. We can adapt this latter interpretation loosely to the social choice framework: if an election is partitioned into two subsets and the winner is determined in each of the subsets, the results should reflect the original winner.
@@ 1. Preliminaries
In this section we define IRV and different types of reinforcement paradoxes. We illustrate these ideas with several examples.
In an election which uses IRV, voters cast preference ballots with a (possibly partial) linear ranking of the candidates. After an election, the ballots are aggregated into a preference profile, which shows the number of each type of ballot cast. For example, Table 1 shows a preference profile involving the three candidates $C$, $G$, and $H$. The number 2211 denotes that 2211 voters ranked $G$ first, $H$ second, and $C$ third. In the real-world IRV elections we analyze in Section 5, voters are allowed to provide partial rankings. Table 1 has three columns where voters do not provide a complete ranking–2003 voters rank only $G$ on their ballots, for example.
An IRV election proceeds in rounds where, in each round, the number of first-place votes for each candidate is calculated. If a candidate has a majority of the (remaining) first-place votes, they are declared the winner. If no candidate receives a majority of votes, the candidate with the fewest first-place votes is eliminated and the names of the remaining candidates on the affected ballots are moved up. The process repeats until a candidate is declared the winner. In the case when partial ballots are allowed, any ballot in which all candidates have been eliminated is considered “exhausted” and plays no further role in the determination of the winner. [^1] In theory there are many ways we could treat exhausted ballots besides removing them from the election. However, as far as we are aware, any actual election office treats exhausted ballots in this fashion.
@@
Initially, the vote totals for Cacy, Giessel, and Holland are 5521, 5653, and 5535, respectively. Cacy has the fewest votes and is eliminated from the election. As a result, 2230 votes are transferred to Giessel and 418 to Holland (the 2873 ballots ranking only Cacy are transferred to no one), and Giessel wins in the last round 7883 votes to 5953.
@@
Table 2 provides a second partition of the ballots, Partition 2, in which $C$ wins under IRV when using only the ballots from $\mathcal{B}_{1}$ and only the ballots from $\mathcal{B}_{2}$, yet when we merge the ballots to recover the original election data, $C$ loses. For both Cacy and Holland, in this election win ++ win == loss in some sense.
@@
Example 2 demonstrates the notion of a reinforcement paradox, the subject of this article. The study of this paradox dates to the work of Young [57]. Young defined a voting method as a function $f$ which takes as input a set of ranked ballots and which outputs an election winner (or a set of winners, if there is some sort of tie). We use a simplified version of Young’s definition which does not address the issue of tied outcomes because ties almost never arise in the elections we study. In particular, none of our real-world elections have a tied outcome in any round. If $\mathcal{B}$ is a set of ballots and $f(\mathcal{B})$ is the winner under voting method $f$, then Young defined the following notion of consistency [57].
@@
A voting method that is not consistent is said to be susceptible to the reinforcement paradox. This paradox has other names, such as the “multiple districts paradox” [22], and is a social choice analogue of the classical Simpson’s Paradox from statistics. In [58], Young showed that consistency and a few other uncontroversial axioms completely characterize positional scoring rules. Thus, any “reasonable” voting method which is not a scoring rule is not consistent. We analyze only IRV and the function $f$ represents that voting method throughout the article.
To investigate the frequency with which reinforcement paradoxes occur, we must first decide what constitutes an “occurrence” or an “instance” of the paradox. There seem to be two viable approaches. The first, used by Courtin et al. [12] and Plassman and Tideman [47], focuses on the ballot sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$. Their work generates ballot sets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ independently using some probability distribution, and the paradox is said to occur if the election winner using $\mathcal{B}_{1}$ is the election winner using $\mathcal{B}_{2}$, but we obtain a different winner for the set $\mathcal{B}_{1}\cup\mathcal{B}_{2}$. A second approach is to start with a ballot set $\mathcal{B}$ and determine if it is possible to decompose into two subsets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ where the winners of $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ match but differ from the winner of the original election. That is, in the second approach we select a single ballot set $\mathcal{B}$ using some probability distribution (or drawing from a set of real-world elections), and say this single election is susceptible to the reinforcement paradox if $\mathcal{B}$ can be partitioned into two sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ such that $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ but that the winner of the two sub-elections is different from $f(\mathcal{B}).$
We do not think either approach is “better” than the other; we view them as complementary. We choose to adopt the second approach for the following reasons. First, we are motivated by a study of real-world elections (see Section 5), and given a single actual election the second approach seems natural. If we focus on decomposing one ballot set into a partition then we can meaningfully ask “Did the 2022 Alaska Special House Election demonstrate susceptibility to a reinforcement paradox?”, and the answer is Yes if we can find a suitable partition of the election’s ballot set.
Second, this approach is in line with notions of “occurrence” for other paradoxes such as monotonicity paradoxes. For example, informally an election is said to demonstrate an upward monotonicity paradox if there exists a set of voters who could turn the winning candidate into a loser by shifting the winner to higher rankings on their ballots [23, 25]. That is, given a single election we demonstrate this paradox by finding a subset of ballots which, when suitably changed, produces a paradoxical outcome. Our approach with reinforcement paradoxes has a similar flavor: while we do not change the ballots, we hunt for subsets of ballots which produce a paradoxical outcome.
Third, much of the interesting work for the first approach has already been covered in [12, 47], but the second approach has not been well-studied. Brandt et al. [5] give results about “minimal” reinforcement paradoxes using the second approach; for example, for a fixed number of candidates and for various voting methods, they give a profile with a minimal number of voters such that the preference profile can be partitioned into two ballots sets which demonstrate the paradox. But otherwise we are unaware of previous results about partitioning a single ballot set. There is much interesting mathematics to explore in the second approach, such as providing necessary and sufficient conditions for an election to admit an appropriate partition of the ballot set. (We do exactly this, for three candidates, in Section 2.) For these reasons, we adopt the following definition for an election to demonstrate a reinforcement paradox.
@@
An election with ballot set $\mathcal{B}$ is susceptible to a reinforcement paradox under a voting method $f$ if there exists a partition of the ballot set $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that $f(\mathcal{B}_{1})=f(\mathcal{B}_{2})$ and $f(\mathcal{B}_{1})\neq f(\mathcal{B}_{1}\cup\mathcal{B}_{2})$.
Under IRV, if an election is susceptible to a reinforcement paradox then the paradox’s dynamics are similar to that of Example 2, in the sense that the elimination order in the two sub-elections are different. It is straightforward to show that if the IRV election using ballots from $\mathcal{B}_{1}$ has the same candidate elimination order as the IRV election using ballots $\mathcal{B}_{2}$ then the winner when using $\mathcal{B}_{1}\cup\mathcal{B}_{2}$ is necessarily the winner of the two sub-elections. Thus, when searching for a partition to demonstrate the paradox with three candidates, we look for sub-elections in which the two candidates in the final round of the election using $\mathcal{B}_{1}$ are not the same as the two candidates in the final round of the election using $\mathcal{B}_{2}$.
When there are three candidates an election can demonstrate a reinforcement paradox in two different ways, leading to the two definitions below. We refer to the candidate with the smallest number of first-place votes as the plurality loser. This candidate can never be the winner in an IRV election since they are eliminated in the first round. We refer to the remaining losing candidate, (who may receive either the largest or second-largest number of first-place votes) as the runner-up.
Definition 3.
An election is susceptible to a plurality loser reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the plurality loser is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ and when using only the ballots in $\mathcal{B}_{2}$.
Definition 4.
An election is susceptible to a runner-up reinforcement paradox if there exists a partition of the ballots $\mathcal{B}=\mathcal{B}_{1}\cup\mathcal{B}_{2}$ such that the runner-up is the IRV winner when using only the ballots in $\mathcal{B}_{1}$ and when using only the ballots in $\mathcal{B}_{2}$.
If an election is susceptible to both a plurality loser and runner-up paradox, we say the election is susceptible to a double reinforcement paradox. We now turn to the problem of determining if a ballot set $\mathcal{B}$ in a three-candidate election can be partitioned to produce a runner-up or plurality loser reinforcement paradox.
2. Necessary and Sufficient Conditions
To determine the likelihood that a three-candidate election is susceptible to a reinforcement paradox, we must be able to characterize when it is possible to partition $\mathcal{B}$ into two sub-elections $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$ such that either the runner-up or the plurality loser wins both sub-elections.
Assume the preference profiles for complete and partial rankings are as shown in Tables 3 and 4 where in each case the expressions $X_{i},Y_{i}$, and $Z_{i}$ indicate the number of voters with the corresponding ranking. Note in Table 4 we do not list ballots of length two, as such ballots convey the same ranking information as the corresponding complete ballots.
@@ 2. Necessary and Sufficient Conditions
1st choice AA AA $B$ $B$ $C$ $C$
2nd choice $B$ $C$ AA $C$ AA $B$
3rd choice $C$ $B$ $C$ AA $B$ AA
@@
1st choice AA AA AA $B$ $B$ $B$ $C$ $C$ $C$
2nd choice $B$ $C$ AA $C$ AA $B$
3rd choice $C$ $B$ $C$ AA $B$ AA
@@
In the complete ballot case, let $X=X_{1}+X_{2}$, $Y=Y_{1}+Y_{2}$, and $Z=Z_{1}+Z_{2}$; in the partial ballot case, let $X^{\prime}=X_{1}+X_{2}+X_{3}$, $Y^{\prime}=Y_{1}+Y_{2}+Y_{3}$, and $Z^{\prime}=Z_{1}+Z_{2}+Z_{3}$. We assume without loss of generality that $A$ is the IRV winner and $C$ is the plurality loser.
First, we observe that if $A$ has a majority of first-place votes in the first round then the election is not susceptible to any kind of reinforcement paradox. That is, we cannot observe a paradox if $X>Y+Z$ (in the complete ballot case) or $X^{\prime}>Y^{\prime}+Z^{\prime}$ (in the partial ballot case). The reason is that if a $A$ has a majority, then they must earn a majority of first-place votes in either $\mathcal{B}_{1}$ or $\mathcal{B}_{2}$ for any partition of the ballot set.
@@
If an election contains a majority candidate then the election is not susceptible to a reinforcement paradox.
Thus, since $A$ is the IRV winner, we assume in the complete ballot case that
$Z<X,\ Z<Y\text{ and }Y+Z_{2}<X+Z_{1}$
and in the partial ballot case that
@@
In addition, as remarked previously, if there is a partition such that either the runner-up $B$ or the plurality loser $C$ wins both sub-elections in two rounds, then the order of elimination in $\mathcal{B}_{1}$ must be different from that in $\mathcal{B}_{2}$. (Suppose, for instance, that $B$ wins both sub-elections in two rounds and that $C$ is the plurality loser in both sub-elections. Using the notation of Table 1, this implies $X^{\prime}+Z_{1}<Y^{\prime}+Z_{2}$ in both sub-elections, which implies the same inequality holds in the original election.) This leads to the following.
Observation 2.
If there is a partition of the electorate in which either the runner-up $B$ or the plurality loser $C$ wins both sub-elections in the second round of counting, then the plurality loser is different in each sub-election.
As the following Propositions indicate, it is also possible for an election to have a partition in which the runner-up wins a majority of first-place votes in one sub-election while winning in round 2 in the other sub-election.